Select and explain a specific ratio of your choice used to analyze financial statements. Explain how to calculate the ratio, what it measures, and what it tells an analyst about the company. Finally, select a company’s recent financial statements, calculate the ratio for your selected company, and explain what this ratio tells you about the company’s financial health or performance.
Answer #1

Solution

When evaluating a stock, investors are always searching for that one golden key measurement that can be obtained by looking at a company's financial statements, but finding a company that ticks off every box is simply not that easy.

To accurately evaluate the financial health and long-term sustainability of a company, a number of financial metrics must be considered. Four main areas of financial health that should be examined are liquidity, solvency, profitability and operating efficiency. However, of the four, likely the best measurement of a company's health is the level of its profitability.

Profitability ratios are a class of financial metrics that are used to assess a business's ability to generate earnings relative to its revenue, operating costs, balance sheet assets, and shareholders' equity over time, using data from a specific point in time. For most profitability ratios, having a higher value relative to a competitor's ratio or relative to the same ratio from a previous period indicates that the company is doing well. Ratios are most informative and useful when used to compare a subject company to other, similar companies, the company's own history, or average ratios for the company's industry as a whole.

There are some common Profitability Ratios, described as follows,

I. Return on Equity: This ratio measures Profitability of equity fund invested the company. It also measures how profitably owner’s funds have been utilized to generate company’s revenues. A high ratio represents better the company is.

Formula: Profit after Tax ÷ Net worth

Where,

Net worth = Equity share capital, and Reserve and Surplus

II. Earnings Per Share: This ratio measures profitability from the point of view of the ordinary shareholder. A high ratio represents better the company is.

Formula: Net Profit ÷ Total no of shares outstanding

III. Dividend Per Share: This ratio measures the amount of dividend distributed by the company to its shareholders. The high ratio represents that the company is having surplus cash.

Formula: Amount Distributed to Shareholders ÷ No of Shares outstanding

IV. Return on Capital Employed: This ratio computes percentage return in the company on the funds invested in the business by its owners. A high ratio represents better the company is.

Formula: Net Operating Profit ÷ Capital Employed × 100

Where,

Capital Employed = Equity share capital, Reserve and Surplus, Debentures and long-term Loans

Or, Total Assets – Current Liability

V. Return on Assets: This ratio measures the earning per $ of assets invested in the company. A high ratio represents better the company is.

Formula: Net Profit ÷ Total Assets

VI. Net Profit Ratio: This ratio measures the overall profitability of company considering all direct as well as indirect cost. A high ratio represents a positive return in the company and better the company is.

Formula: Net Profit ÷ Total Revenues × 100

Where,

Net Profit = Gross Profit + Indirect Income – Indirect Expenses

Now for analysis purpose, we have taken the Financial Statement of CVS Health which is a renowned healthcare service provider in United States for the year ending on December 31st, 2018. Link for the full annual report and consolidated financial report part given below,

Complete Annual Report: https://drive.google.com/file/d/1wKi_znhfRTHZ6Um6-ZgTvJ0lncSUh0Vt/view?usp=sharing

Consolidated Financial Statement part: https://drive.google.com/file/d/1PLOBxi64UQn2fek9L8bfoy-f9b71AV1D/view?usp=sharing

Now on the basis of the information, we will measure the financial performance as follows,

Ratios

2018

2017

Notes

Return on Equity

(-) 594 / 58225 = (-) 1.02% (Approx.)

6622 / 37691 = 17.569% (Approx.)

Refer to the Consolidated FS,

  1. Profit After Tax = Net Income / (Loss) attributable
  2. Net Worth = Total CVS Health Shareholders’ Equity

Comment: As it measures how profitably owner’s funds have been utilized to generate company’s revenues, it can be concluded there was some mismanagement due to which the company was not able to utilize the revenues properly this year than the previous year, resulting to which the company’s ROE has been massively reduced.

Earnings Per Share ($)

(-) 0.57

6.47

Refer to the Consolidated FS, the basic EPS has been considered here.

Comment: As it measures profitability from the point of view of the ordinary shareholder, it can be concluded that the company is not performing well, and due to that the ordinary shareholders may lose their interest in such investment.

Dividend Per Share ($)

2.00

2.00

Clearly mentioned in the Consolidated FS.

Return on Capital Employed

4021 / 152447 = 2.638% (Approx.)

9538 / 64483 = 14.791% (Approx.)

Refer to the Consolidated FS,

  1. Net Operating Profit = Operating Income
  2. Capital Employed = Total Assets – Total Current Liabilities
  1. For 2018 = 196456 – 44009 = 152447
  2. For 2017 = 95131 – 30648 = 64483

Comment: As this ratio computes percentage return in the company on the funds invested in the business by its owners, it can be said that due to company’s mismanagement, percentage return has been fall massively in current year.

Return on Assets

(-) 594 / 196456 = 0.302% (Approx.)

6622 / 95131 = 6.961% (Approx.)

Refer to the Consolidated FS,

  1. Net Profit = Net Income / (Loss) attributable
  2. Total Assets = Given in FS

Comment: As this ratio measures the earning per $ of assets invested in the company, it can be seen that the per $ earning has been reduced massively.

Net Profit Ratio

(-) 596 / 194579 = (-) 0.306% (Approx.)

6622 / 184786 = 3.584% (Approx.)

Refer to the Consolidated FS,

  1. Net Profit = Net Income / (Loss) attributable
  2. Total Revenues = Mentioned in FS

Comment: As this ratio measures the overall profitability of company considering all direct as well as indirect cost, it can be concluded that the overall profitability has been affected adversely in current year.

Overall Comment: As said previously, profitability ratios are a class of financial metrics that are used to assess a business's ability to generate earnings. On the basis of the calculations made above, the company has a negative profitability in current year than previous year. Hence it can be concluded that the financial performance of the company in “Profitability” aspect has become quite bad that previous year.